# Local Intrinsic Dimensionality of Ground Motion Data for Early Detection of Complex Catastrophic Slope Failure

### v1

@@ 3.4. Spatiotemporal LID
(10) $st\text{-LID}({\bm{x}}_{t})=\sigma(s\text{-LID}({\bm{x}}_{t}))\times\sigma(t\text{-LID}({\bm{x}}_{t})).$
@@
3.5. Real-World Failure Detection via $st$-LID
@@ 4.1. Datasets
Figure 5. Visualization of the ground truth in 4 datasets (columns) and corresponding detection results from 5 methods (rows). Detection results are colored based on the likelihood of the failure (see color-bar for detailed range and order)
Table 1. Precision Score $Prec.$ ($\uparrow$) & Lead-time $\Delta t$ ($\uparrow$) results from all methods. $\uparrow$ represents higher values are better. Best performance are highlighted in bold format.
@@ 4.4. Computation Time
Table 2. Computation time of each failure detection methods.
@@
4.5. Real-World Failure Detection Simulation
@@
Figure 6. Simulation of real-world failure detection using stst-LID. The monitored point with the highest $st$-LID is marked by $\star$.
@@ 5. Conclusion
Appendix A Displacement Time Series Data
@@
Figure 7. Visualization of the displacement time series.

### v2

@@ 3.4 Spatiotemporal LID
$st\text{-LID}({\bm{x}}_{t})=\sigma(s\text{-LID}({\bm{x}}_{t}))\times\sigma(t\text{-LID}({\bm{x}}_{t})).$ (10)
@@
3.5 Real-World Failure Detection via $st$-LID
@@ 4.1 Datasets
Figure 5: Visualization of the ground truth in 4 datasets (columns) and corresponding detection results from 5 methods (rows). Detection results are colored based on the likelihood of the failure (see color-bar for detailed range and order)
Table 1: Precision Score $Prec.$ ($\uparrow$) & Lead-time $\Delta t$ ($\uparrow$) results from all methods. $\uparrow$ represents higher values are better. Best performance are highlighted in bold format.
@@ 4.4 Computation Time
Table 2: Computation time of each failure detection methods.
@@
4.5 Real-World Failure Detection Simulation
@@
Figure 6: Simulation of real-world failure detection using stst-LID. The monitored point with the highest $st$-LID is marked by $\star$.
@@ 5 Conclusion
6 Acknowledgment
We would like to thank the GroundProbe for providing the datasets and support for this study.
Appendix 0.A Displacement Time Series Data
@@
Figure 7: Visualization of the displacement time series.